Contribution Margin and Contribution Margin Ratio

For a recent year, McDugal's company-owned restaurants had the following sales and expenses (in millions):

Sales   $36,800
Food and packaging   $9,672
Payroll 9,800
Occupancy (rent, depreciation, etc.) 11,728
General, selling, and admin. expenses    5,600
Other expense 740
Total expenses (37,540)
Operating income (loss) $(740)

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What would have been the operating income or loss for the recent year if sales had been $2,200 million more?
$ ______ million

b. To achieve break even for the recent year, by how much would sales need to increase? Enter your anwer in million rounded to the nearest whole number.
$ _______ million
Answer #1

Variable cost = 9672+9800+(5600*40%) = 21712

Contribution margin ratio = (36800-21712)/36800 = 41%

a) Net operating income = (2200*41%-740) = 162 Million

b) Break even sales = 15828/.41 = 38605

Increase = 38605-36800 = 1805 Million